OpenStudy (anonymous):

HEEELP! How High Is It? How do you get the 66 part?

OpenStudy (anonymous):



OpenStudy (anonymous):

The cylinder has a surface area of 660 cm2. What is the height of the cylinder?

OpenStudy (anonymous):

\[SA=2\pi r^2+2\pi rh\] \[SA=2\pi 7^2+2\pi 7h\] \[660=98\pi +14\pi h\] \[{h}={66\[\therefore h \approx 8.01\]0-98\pi} \over 14\pi\]

OpenStudy (anonymous):

h = 8.01 approximately

OpenStudy (anonymous):

how did you get the 66 and the 8.01

OpenStudy (anonymous):

?